# Bose-Einstein condensation of THz photons in an optical microcavity with Landau-quantized electrons

### v1

@@ I Introduction
When decreasing the temperature in a system of Bose particles, the associated thermal de-Broglie wavelength increases and, eventually, exceeds the mean interparticle distance. That is where quantum effects manifest themselves, leading to Bose-Einstein condensation (BEC) as observed experimentally for cold atomic gases [1, 2, 3]. Further progress in experimental techniques led to observation of condensation signatures in quasiparticle systems such as excitons [4, 5], excitonic polaritons [6, 7] and microcavity photons [8]. Due to finite quasiparticle lifetime, these are inherently open driven-dissipative systems, which encourages using them as a playground for investigating non-equilibrium phase transitions, critical behaviour, etc. In addition, polaritonic/photonic condensates are of special interest due to facilitated access to the optical emission, which readily implies possible applications as coherent laser-type sources [9].
The main objective of the current work is considering a THz radiation source based on the emission of a microcavity photon BEC. In order to achieve the desired condensed state one needs to provide some inter-photon interaction mechanism, i.e. introduce some ”active medium” , which acts as a source of the essential nonlinearity [10]. Dye molecules are an example in the renowned experiment by the group of J. Klaers [8]. In our setup we propose utilizing a 2D electron gas (2DEG) in magnetic field as a tunable photonic interaction mediator.
@@
Currently, one has a bunch of techniques to choose from for THz generation. Namely, free electron lasers, quantum cascade lasers, optical laser based nonlinear converters, etc. [11, 16, 19] We claim that microcavity photonic condensate also belongs to this list, being a decent radiation source.
The article is organized as follows. We open Sec. II with a presentation of the principal idea for THz photon condensation and a schematic of the coherent light source, which is followed by a microscopic model. Then follows the core part of the paper: in Sec. III we study the details of kinetic processes which feed and deplete the photonic condensate. The subsequent Sec. IV is devoted to assessing and optimizing device performance by properly choosing control parameters (magnetic field, cavity Q factor, etc.) in terms of the output power and efficiency. In addition, we address several technical issues of creating a device for practical usage. Sec. V summarizes the results of our study. Finally, Appendices A-F present calculation details which, for the sake of clarity, are omitted in the main text.
@@ II The THz coherent source
II.1 The principal idea
The scheme of the proposed device is presented in Fig. 1. The idea is to create conditions for Bose-Einstein condensation in a single photonic mode of a microcavity illuminated by a thermal light source in order for the emission for this mode outside the cavity to be coherent.
@@ II.1 The principal idea
The main challenge in designing a device of the type described above is the dissipation. Photonic modes of the microcavity constitute an open system constantly exposed to interaction with the lattice and external pumping source. Due to the former, photons transmit the absorbed energy to the semiconductor, thus equilibrating with the lattice at relatively low temperatures. For an occupation buildup in a single mode, one has to increase the pumping rate of that mode and/or to suppress the leakage out of it. We propose using electron cyclotron resonance (ECR) to control the photon absorption rate by the 2DEG. As will be carefully demonstrated further, for a 2DEG in magnetic field, the rate of photonic outflow from a single optical mode of the microcavity (due to absorption by the electron gas) depends on its frequency in a periodic manner. This effect has been experimentally investigated for heterostructures in THz range (e.g. see [20]). As depicted in Fig. 2, peak values are close to the resonant absorption frequencies (the shapes of the peaks are due to the details of electron-phonon interaction). In contrast, the modes in the regions of low leakage are effectively isolated from the lattice.
@@
However, low transmittance of the mirrors also prevents the essential efficient pumping of the condensate mode by the black body. To overcome this obstacle, we note that quantized electrons well-mediate interaction between photons in modes spaced by twice the cyclotron frequency $\omega_{H}$ (see (7)). Having some photonic modes (hereafter ”pumping modes” ) with small Q-factor (thus, penetrable for external pumping) in the next, but one minimum of the leakage rate profile provides the necessary source of influx to the condensate mode. In addition, to broaden the frequency region of pumping modes, as shown in Fig. 2, we propose introducing a magnetic field gradient over the semiconductor sample.
Note that besides the condensate mode with zero in-plane momentum, there are multiple modes of the same energy (hereafter ”satellite modes”), which are also well-coupled to the condensate mode via elastic photon scattering on electrons.
Altogether, the pumping modes along with the condensate mode with its satellites make up a photonic subsystem, which is effectively isolated from both the lattice and the other photonic modes and exposed to black body pumping and dissipation through the Bragg mirrors.
Figure 2: Typical dependence of the photon absorption rate by the 2DEG $\nu_{\rm loss}$ as a function of photonic mode frequency (left axis, red dashed graph). The high-Q condensate mode along with a bunch of low Q pumping modes are depicted in blue.
@@ II.2 Microscopic model
As a microscopic model of the described device, we use the following Hamiltonian, which describes all the essential subsystems and interactions between them:
$\qquad\qquad\qquad\hat{H}=\underbrace{\sum_{\bm{q},\xi}\hbar\omega_{\bm{q}}\hat{c}^{\dagger}_{\bm{q},\xi}\hat{c}_{\bm{q},\xi}}_{\rm microcavity\;photons}+\underbrace{\sum_{\bm{k},\zeta}E_{\rm ph}({\bm{k}},\zeta)\hat{b}^{\dagger}_{\bm{k},\zeta}\hat{b}_{\bm{k},\zeta}}_{\rm acoustic\;and\;optical\;phonons}+\underbrace{\sum_{\bm{p},s}E_{\bm{p}}^{s}\hat{a}_{\bm{p},s}^{\dagger}\hat{a}_{\bm{p},s}}_{\rm Landau-quantized\;electrons}$
${+}\underbrace{\sum_{\begin{subarray}{c}{\bm{p},\bm{p}_{1},s}\\ {\bm{k},\zeta}\end{subarray}}\kappa_{\bm{p}\bm{p}_{1}}(\bm{k},\zeta)\left(\hat{b}_{\bm{k},\zeta}{+}\hat{b}_{\bm{k},\zeta}^{\dagger}\right)\hat{a}_{\bm{p}_{1},s}^{\dagger}\hat{a}_{\bm{p},s}}_{\rm electron-phonon\;interaction}{+}\underbrace{\sum_{\begin{subarray}{c}\bm{p}\bm{p}_{1}\\ \bm{q},\xi\end{subarray}}\left(\lambda_{\bm{p}\bm{p}_{1}}^{-}(\bm{q},\xi)\hat{c}_{\bm{q},\xi}+\lambda_{\bm{p}\bm{p}_{1}}^{+}(\bm{q},{\shortminus\xi})\hat{c}_{\bm{q},{\shortminus\xi}}^{\dagger}\right)\hat{a}_{\bm{p}_{1},\xi}^{\dagger}\hat{a}_{\bm{p},{\shortminus}\xi}}_{\rm electron-photon\;interaction}.$ (1)
Here $\bm{q}=\bm{q}_{\parallel}+q_{z}(l)\bm{e}_{z}$ is the photon 3D momentum with $\bm{q}_{\parallel}=q_{x}\bm{e}_{x}+q_{y}\bm{e}_{y}$ being the in-plane momentum. The transverse quantization level number is denoted by $l$. Hence, a momentum summation is over both the 2D-momentum and transverse level, $\sum_{\bm{q}}=\sum_{\bm{q}_{\parallel},l}$. The $\hat{c}_{\bm{q},\xi}$ is the photon annihilation operator in a state with the momentum $\bm{q}$ and polarization $\xi$ ($\xi=\pm 1$ for right/left circular polarization). Thus, the frequency is given as
@@
where $\varepsilon$ is the dielectric constant of the material of the inside of the microcavity and $c$ is the vacuum speed of light. The condensate photon frequency is denoted by $\omega_{c}\equiv\omega_{\bm{0}}=cq_{z}(0)/\sqrt{\varepsilon}$.
The same decomposition is adopted for photon momentum $\bm{k}=\bm{k}_{\parallel}+k_{z}\bm{e}_{z}$ with $\bm{k}_{\parallel}=k_{x}\bm{e}_{x}+k_{y}\bm{e}_{y}$. The operator $\hat{b}_{\bm{k},\zeta}$ annihilates a phonon with ${\bm{k}}$ momentum in a mode $\zeta$ (acoustic and optical modes). Dispersions are given by $E_{\rm acoustic}(k)=ku$ and $E_{\rm optical}={\rm const}$, where $u$ is the speed of sound and $k=\sqrt{{\bm{k}}_{\parallel}^{2}+k_{z}^{2}}$.
@@
We use a collective index for the electron state: $\bm{p}\equiv\{p_{x},n,\mathfrak{L}\}$: $p_{x}$ is the momentum projection on in-plane direction $Ox$, $n$ stands for the Landau level number, $\mathfrak{L}$ is the transverse quantization level number. Thus, $\hat{a}_{\bm{p},s}$ annihilates a photon from trapping potential level $\mathfrak{L}$, Landau level $n$ with spin projection $s$. Note, that the electron spectrum does not depend on the $p_{x}$, thus we will further eliminate it.
@@
In addition, in subsequent sections we consider magnetic field with spatial gradient in order to facilitate energy absorption from the thermal source to the pumping modes. As estimated in Appendix B, the gradient can be chosen weak enough not to affect the Hamiltonian, we just treat $\omega_{H}$ above as the average magnetic field across the area of the device.
@@
$\lambda_{\bm{p}\bm{p}_{1}}^{\pm}(\bm{q},\xi)\equiv\lambda_{nm}^{\pm}(\bm{q},\xi)$
$\kappa_{\bm{p}\bm{p}_{1}}(\bm{k},\zeta)\equiv\kappa_{nm}(\bm{k},\zeta).$
@@ III Kinetics
Figure 3: A schematic of all the relevant processes in the photonic subsystem of interest. (a) Dotted horizontal lines on the $\omega(\bm{q}_{\parallel})$ plane depict modes of the same transverse momentum. Relevant photonic modes are depicted in blue. Red arrows show particle exchange processes. The black wavy line depicts a phonon, which is emitted when 2DEG absorbs a condensate photon. (b) Thermal radiation intensity frequency profile with the pumping region being shaded.
@@ III.1 Condensate mode
In order to describe the kinetic processes, which feed/deplete condensate mode, we make use of the Keldysh technique following  [22]. Rates of photonic exchange are given in terms of the retarded self-energy term $\Sigma^{R}(t-t^{\prime})=\sigma^{R}\delta(t-t^{\prime})$:
$\frac{\partial N_{0}}{\partial t}=N_{0}\Im[\sigma^{R}]=\!\!\!\!\sum_{j\in\Big{\{}\begin{subarray}{c}{\rm pump}\\ {\rm loss}\\ {\rm satell}\end{subarray}\Big{\}}}\left.\!\!\!\!\!\!\partial_{t}N_{0}\right|_{j}.$ (4)
The index $j$ runs over all the possible photon scattering channels. For each of them we may express the corresponding term in a form similar to the Fermi’s golden rule:
$\frac{1}{N_{0}}\partial_{t}N_{0}\big{|}_{j}=\frac{2\pi}{\hbar}\sum_{\bm{p},\bm{p}^{\prime}}|A^{(j)}_{\bm{p},\bm{p}^{\prime}}|^{2}\delta\left(E^{(j)}_{\text{initial}}(\bm{p})-E^{(j)}_{\text{final}}(\bm{p}^{\prime})\right)$
×${\times}\!\left[N^{(j\,in)}_{\text{init}}(\bm{p})N^{(j\,in)}_{\text{final}}(\bm{p}^{\prime}){-}N^{(j\,out)}_{\text{init}}(\bm{p})N^{(j\,out)}_{\text{final}}(\bm{p}^{\prime})\right].$ (5)
Here we took advantage of macroscopic condensate occupation $N_{0}\gg 1$ to consider solely the stimulated processes. with corresponding amplitudes $A^{(j)}_{\bm{p},\bm{p}^{\prime}}$, sum of energies $E^{(j)}_{\rm initial/final}$ and occupation number products $N^{(j\,in/out)}_{\rm init/final}$ of initial and final scattering states.
@@
(i) Condensate replenishment from the pumping modes
Particle exchange between the condensate and pumping modes is due to resonant scattering on electrons. The corresponding Feynman diagrams are given in Fig. 4:
@@
$\partial_{t}\ln(N_{0})\Big{|}_{\rm pump}\!\!\!=\nu_{\rm in}N_{\rm pump}^{\rm ph}-\nu_{\rm out}(1+N_{\rm pump}^{\rm ph}),$
which can be recast as follows:
$\partial_{t}\ln(N_{0})\Big{|}_{\rm pump}\!\!\!=N_{{\rm pump}}^{\rm ph}\nu_{\rm pump}-\nu_{\rm reverse}.$ (6)
Here $N_{{\rm pump}}^{\rm ph}$ stands for the occupation of the pumping modes, $\nu_{\rm pump}=\nu_{\rm in}-\nu_{\rm out}$, $\nu_{\rm reverse}=\nu_{\rm out}$:
$\nu_{\rm pump}{=}\frac{2\pi}{\hbar}\!\!\!\sum_{\begin{subarray}{c}\bm{p},{\bm{q}}_{\rm pump}\\ {s=\pm 1}\end{subarray}}\!\!\left|\frac{\lambda_{n+1\shortminus s,n+1}^{-}(\bm{q}_{\rm pump})\lambda^{+}_{n+1,n+1+s}(\bm{0})}{E^{s}_{n+1+s}+s\hbar\omega_{c}-E_{n+1}^{-s}}\right|^{2}$
×$\times\left(N^{\rm el}_{n,s}{-}N^{\rm el}_{n+2,s}\right)\delta(\hbar\omega_{\rm pump}{+}E_{n}^{s}{-}\hbar\omega_{c}{-}E_{n+2}^{s}),$ (7)
$\nu_{\rm reverse}{=}\frac{2\pi}{\hbar}\!\!\sum_{\begin{subarray}{c}\bm{p},\bm{q}_{\rm pump}\\ {s=\pm 1}\end{subarray}}\!\left|\frac{\lambda^{-}_{n+1+s,n+1}(\bm{0})\lambda_{n+1,n+1\shortminus s}^{+}(\bm{q}_{\rm pump})}{E^{s}_{n+1+s}+s\hbar\omega_{c}-E_{n+1}^{-s}}\right|^{2}$
×$\times N^{\rm el}_{n+2,s}\left(1{-}N^{\rm el}_{n,s}\right)\delta(\hbar\omega_{\rm pump}{+}E_{n}^{s}{-}\hbar\omega_{c}{-}E_{n+2}^{s}).$ (8)
The 3D momentum of the pumping photon is denoted by $\bm{q}_{\rm pump}$, electron level occupation is introduced as
@@
The diagrams in Fig. 5 describe photon absorption by 2DEG with phonon emission.
@@
The rate of this process is given by the following expression:
$\nu_{\rm loss}{=}\frac{2\pi}{\hbar}\!\!\sum_{\bm{p},\bm{k},\zeta}\!\left|\!\frac{\lambda^{-}_{n,n\shortminus 1}(\bm{0})\kappa_{n\shortminus 1,m}(\bm{k},\zeta)}{E^{+}_{n}{+}\hbar\omega_{c}{-}E_{n\shortminus 1}^{-}}{+}\frac{\kappa_{n,m+1}(\bm{k},\zeta)\lambda^{-}_{m+1,m}(\bm{0})}{E^{-}_{m}{-}\hbar\omega_{c}{-}E_{m+1}^{+}}\!\right|^{2}\!\!\!\frac{N^{\rm el}_{n,+1}{-}N^{\rm el}_{m,\shortminus 1}}{e^{E_{\rm ph}(\bm{k},\zeta)/T}-1}\delta(\hbar\omega_{c}{+}E_{n}^{+}{-}E_{m}^{-}{-}E_{\rm ph}({\bm{k}},\zeta)).$ (9)
@@
for optical ones. The quantities $\Delta n_{a/{\rm opt}}$ denote the number of Landau levels between the scattering states of the electron.

### v2

@@ I Introduction
When decreasing the temperature in a system of Bose particles, the associated thermal de-Broglie wavelength increases and, eventually, exceeds the mean interparticle distance. That is where quantum effects manifest themselves, leading to Bose-Einstein condensation (BEC) as observed experimentally for cold atomic gases [1, 2, 3]. Further progress in experimental techniques led to observation of condensation signatures in quasiparticle systems such as excitons [4, 5], magnons [6], excitonic polaritons [7, 8] and microcavity photons [9]. Due to finite quasiparticle lifetime, these are inherently open driven-dissipative systems, which encourages using them as a playground for investigating non-equilibrium phase transitions, critical behaviour, etc. In addition, polaritonic/photonic condensates are of special interest due to facilitated access to the optical emission, which readily implies possible applications as coherent laser-type sources [10].
For the occupation build-up in a single mode in these quasiparticle systems, one has to provide a source of enhanced influx to compensate the losses or minimize these losses directly. Typically, particles in the condensate are generated through scattering from high-energy states (as observed in exciton, exciton-polariton, and magnon condensates) or with the assistance of a gain medium. Our focus in this paper is on the photonic condensate in an optical microcavity. In these setups, the relaxation of the excited gain medium is routinely employed, given the negligible photon interaction. Experimental studies have employed dye molecules [9, 11, 12] or semiconductor layers in VCSELs [13, 14, 15] as the gain medium. The kinetics of condensate interaction with the gain medium in these experimental setups has been also investigated extensively [16, 17, 18, 19, 20, 21, 22] (see also the review [23]).
The condensate frequency in the mentioned experiments is in visible/IR range. Generating condensate of THz frequency in the same manner is not straightforward. Besides the need to find a medium with low excitation energy, it would require reduced temperatures (note, though, that coherent THz generation was realized with a coupled-cavity scheme with infrared VCSELs [24]).
We propose and analyze an experimental scheme, which has a potential to overcome these difficulties. In terms of the geometry it is a typical VSCEL-type device, although the operation mechanism is different. Namely, we consider a photonic microcavity with an n-doped semiconductor layer with its electrons quantized by intense magnetic field. This setup allows us to suppress losses from the cavity’s condensate mode via the emission of lattice photons. Additionally, our design provides conditions for the efficient relaxation of high-energy photons, injected by a thermal light source, into the condensate mode. In a sense, we present a means to adapt the kinetic condensation mechanism (typically observed for excitons and exciton-polaritons) for non-interacting photons within the microcavity. Thus, our theoretical analysis is qualitatively different from that present for existing experimental setups for photonic condensation.
@@
Currently, one has a bunch of techniques to choose from for THz generation. Namely, free electron lasers, quantum cascade lasers, optical laser based nonlinear converters, etc. [25, 30, 33]. We claim that microcavity photonic condensate also belongs to this list, being a decent radiation source.
The article is organized as follows. We open Sec. II with a presentation of the underlying principle of THz photon condensation and a schematic of the coherent light source, which is followed by a microscopic model. Then follows the core part of the paper: in Sec. III we study the details of kinetic processes which feed and deplete the photonic condensate. The subsequent Sec. IV is devoted to assessing and optimizing device performance by properly choosing control parameters (magnetic field, cavity Q factor, etc.) in terms of the output power and efficiency. In addition, we address several technical issues of creating a device for practical usage. Sec. V summarizes the results of our study. Finally, Appendices A-F present calculation details which, for the sake of clarity, are omitted in the main text.
@@ II The THz coherent source
II.1 The underlying principle
The scheme of the proposed device is presented in Fig. 1. The aim is to achieve macroscopic population of a single photonic mode in an optical microcavity, illuminated by a thermal light source, in order for the emission for this mode outside the cavity to be coherent.
@@ II.1 The underlying principle
The main challenge in designing a device of the type described above is the dissipation. Photonic modes of the microcavity constitute an open system constantly exposed to interaction with the lattice and external pumping source. Due to the former, photons transmit the absorbed energy to the semiconductor, thus equilibrating with the lattice at relatively low temperatures. For an occupation buildup in a single mode, one has to increase the pumping rate of that mode and to suppress the leakage out of it. We propose using electron cyclotron resonance (ECR) to control the photon absorption rate by the 2DEG. As will be carefully demonstrated further, for a 2DEG in magnetic field, the rate of photonic outflow from a single optical mode of the microcavity (due to absorption by the electron gas) depends on its frequency in a periodic manner. This effect has been experimentally investigated for heterostructures in THz range (e.g. see [34]). As schematically depicted in Fig. 2, peak values are close to the resonant absorption frequencies (the shapes of the peaks are due to the details of electron-phonon interaction). In contrast, the modes in the regions of low leakage are effectively isolated from the lattice.
@@
However, low transmittance of the mirrors also prevents the essential efficient pumping of the condensate mode by the black body. To overcome this obstacle, we note that quantized electrons well-mediate interaction between photons in modes spaced by twice the cyclotron frequency $\omega_{H}$ (as will be demonstrated further, see (7)). Having some photonic modes with a small $Q$-factor (thus, penetrable for external pumping) in the next, but one minimum of the leakage rate profile provides the necessary source of influx to the condensate mode. In addition, to broaden the frequency range $\Delta\omega_{\rm res}$ of these modes, as shown in Fig. 2, we propose introducing a magnetic field gradient over the semiconductor sample. We will further refer to these modes collectively as ”pumping reservoir”.
Note that, besides the condensate mode with zero in-plane momentum, there are multiple modes of the same energy (hereafter ”satellite modes”, see Fig. 3), which are also well-coupled to the condensate mode via elastic photon scattering on electrons.
Altogether, the pumping reservoir along with the condensate mode with its satellites make up a photonic subsystem, which is effectively isolated from both the lattice and the other photonic modes and exposed to black body pumping and dissipation through the Bragg mirrors. The kinetic processes in this subsystem are of interest in this work.
Figure 2: Typical dependence of the photon absorption rate by the 2DEG $\varkappa_{\rm loss}$ as a function of photonic mode frequency (left axis, red dashed graph). The high-$Q$ condensate mode along with a bunch of low $Q$ pumping modes are depicted in blue.
@@ II.2 Microscopic model
To model the described device microscopically, we employ the following Hamiltonian, which characterizes all essential subsystems and their interactions:
$\hat{H}=\underbrace{\sum_{\bm{q},\xi}\hbar\omega_{\bm{q}}\hat{c}^{\dagger}_{\bm{q},\xi}\hat{c}_{\bm{q},\xi}}_{\rm microcavity\;photons}+\underbrace{\sum_{\bm{k},\zeta}E_{\rm ph}({\bm{k}},\zeta)\hat{b}^{\dagger}_{\bm{k},\zeta}\hat{b}_{\bm{k},\zeta}}_{\rm acoustic\;and\;optical\;phonons}$
$+\underbrace{\sum_{\bm{p},s}E_{\bm{p}}^{s}\hat{a}_{\bm{p},s}^{\dagger}\hat{a}_{\bm{p},s}}_{{\begin{subarray}{c}\text{Landau-quantized}\\ \text{electrons}\end{subarray}}}\!\!{+}\underbrace{\sum_{\begin{subarray}{c}{\bm{p},\bm{p}_{1},s}\\ {\bm{k},\zeta}\end{subarray}}\kappa_{\bm{p}\bm{p}_{1}}(\bm{k},\zeta)\left(\hat{b}_{\bm{k},\zeta}{+}\hat{b}_{\bm{k},\zeta}^{\dagger}\right)\hat{a}_{\bm{p}_{1},s}^{\dagger}\hat{a}_{\bm{p},s}}_{\rm electron-phonon\;interaction}$
${+}\underbrace{\sum_{\begin{subarray}{c}\bm{p}\bm{p}_{1}\\ \bm{q},\xi\end{subarray}}\left(\lambda_{\bm{p}\bm{p}_{1}}^{-}(\bm{q},\xi)\hat{c}_{\bm{q},\xi}+\lambda_{\bm{p}\bm{p}_{1}}^{+}(\bm{q},{\shortminus\xi})\hat{c}_{\bm{q},{\shortminus\xi}}^{\dagger}\right)\hat{a}_{\bm{p}_{1},\xi}^{\dagger}\hat{a}_{\bm{p},{\shortminus}\xi}}_{\rm electron-photon\;interaction}.$ (1)
Here $\bm{q}=\bm{q}_{\parallel}+q_{z}(l)\bm{e}_{z}$ is the photon 3D momentum with $\bm{q}_{\parallel}=q_{x}\bm{e}_{x}+q_{y}\bm{e}_{y}$ being the in-plane momentum. The transverse quantization level number is denoted by $l$. Hence, a momentum summation is over both the 2D-momentum and the transverse level, $\sum_{\bm{q}}=\sum_{\bm{q}_{\parallel},l}$. The $\hat{c}_{\bm{q},\xi}$ is the photon annihilation operator in a state with the momentum $\bm{q}$ and polarization $\xi$ ($\xi=\pm 1$ for right/left circular polarization). Thus, the frequency is given as
@@
where $\varepsilon$ is the dielectric constant of the medium inside the microcavity and $c$ is the vacuum speed of light. The condensate photon frequency is denoted by $\omega_{c}\equiv\omega_{\bm{0}}=cq_{z}(0)/\sqrt{\varepsilon}$.
The same decomposition is adopted for the phonon momentum $\bm{k}=\bm{k}_{\parallel}+k_{z}\bm{e}_{z}$ with $\bm{k}_{\parallel}=k_{x}\bm{e}_{x}+k_{y}\bm{e}_{y}$. The operator $\hat{b}_{\bm{k},\zeta}$ annihilates a phonon with $\bm{k}$} momentum in a mode $\zeta$ (acoustic and optical modes). Dispersions are given by $E_{\rm acoustic}(k)=ku$ and $E_{\rm optical}={\rm const}$, where $u$ is the speed of sound and $k=\sqrt{{\bm{k}}_{\parallel}^{2}+k_{z}^{2}}$.
@@
We use a collective index for the electron state: $\bm{p}_{n}\equiv\{p_{x},n,\mathfrak{L}\}$: $p_{x}$ is the momentum projection on in-plane direction $Ox$, $n$ stands for the Landau level number, $\mathfrak{L}$ is the transverse quantization level number. Thus, $\hat{a}_{\bm{p},s}$ annihilates a photon from trapping potential level $\mathfrak{L}$, Landau level $n$ with spin projection $s$. Note, that the electron spectrum does not depend on the $p_{x}$, thus we will further eliminate it.
@@
As already mentioned, in subsequent sections we consider magnetic field with spatial gradient in order to facilitate energy absorption from the thermal source to the pumping modes. As estimated in Appendix B, the gradient can be chosen weak enough not to affect the Hamiltonian, we just treat $\omega_{H}$ above as the average magnetic field across the area of the device.
@@
$\lambda_{\bm{p}\bm{p}_{1}}^{\pm}(\bm{q},\xi)\equiv\lambda_{nm}^{\pm}(\bm{q},\xi)\delta_{p_{x},p_{1x}}$
$\kappa_{\bm{p}\bm{p}_{1}}(\bm{k},\zeta)\equiv\kappa_{nm}(\bm{k},\zeta)\delta_{p_{x},p_{1x}}.$
@@ III Kinetics
Figure 3: A schematic of all the relevant processes in the photonic subsystem of interest (blue shaded regions). (a) Dotted lines on the $\omega(\bm{q}_{\parallel})$ plane depict modes of the same transverse momentum(quantized). Red arrows depict particle exchange processes, with their orientation defining the positive sign convention. Double arrows highlight the dominant processes in the condensate state. The black wavy line depicts a phonon, which is emitted when 2DEG absorbs a photon. (b) Thermal radiation intensity frequency profile with the pumping region being shaded.
@@ III.1 Condensate mode
In order to describe the kinetic processes, which feed/deplete condensate mode, we make use of the Keldysh technique  [36]. Rates of photonic exchange are given in terms of the retarded self-energy term $\Sigma^{R}(t-t^{\prime})=\sigma^{R}\delta(t-t^{\prime})$:
$\frac{\partial N_{0}}{\partial t}=\Im[\sigma^{R}]=\!\!\!\!\sum_{j\in\Big{\{}\begin{subarray}{c}{\rm pump}\\ {\rm loss}\\ {\rm sat}\end{subarray}\Big{\}}}\left.\!\!\!\!\!\!\!d_{t}N_{0}\right|_{j}.$ (4)
The index $j$ runs over all the possible photon scattering channels. For each of them we may express the corresponding term in the Fermi’s golden rule form:
$\ \ \ d_{t}N_{0}\big{|}_{j}=\frac{2\pi}{\hbar}\sum_{\bm{p},\bm{p}^{\prime}}|A^{(j)}_{\bm{p},\bm{p}^{\prime}}|^{2}\delta\left(E^{(j)}_{\text{initial}}(\bm{p})-E^{(j)}_{\text{final}}(\bm{p}^{\prime})\right){\times}$
$\!\!\!\!\left[N^{(j\,in)}_{\text{init}}(\bm{p})N^{(j\,in)}_{\text{final}}(\bm{p}^{\prime})(1{+}N_{0}){-}N^{(j\,out)}_{\text{init}}(\bm{p})N^{(j\,out)}_{\text{final}}(\bm{p}^{\prime})N_{0}\right].$ (5)
Here we used the amplitudes $A^{(j)}_{\bm{p},\bm{p}^{\prime}}$, total energies $E^{(j)}_{\rm initial/final}$ and occupation number products $N^{(j\,in/out)}_{\rm init/final}$ of initial and final scattering states.
@@
(i) Condensate replenishment from the pumping reservoir
Particle exchange between the condensate and pumping reservoir is due to resonant scattering on electrons. The corresponding Feynman diagrams are given in Fig. 4:
@@
$d_{t}N_{0}\Big{|}_{\rm pump}\!\!\!=RN_{\rm res}(1+N_{0})-r(1+N_{\rm res})N_{0},$ (6)
In this equation and further we treat all the pumping modes collectively as a single reservoir with $N_{\rm res}$ being its occupation. By doing so, we assume that the pumping range broadening due to the magnetic field gradient is narrow enough for all the kinetic rates and the black body photon occupation not to change significantly over this interval.
The rates introduced above are given as follows:
$R{=}\frac{2\pi}{\hbar}\!\!\!\sum_{\begin{subarray}{c}\bm{p}_{n},{\bm{q}}_{\rm res}\\ {s=\pm 1}\end{subarray}}\!\!\left|\frac{\lambda_{n+1\shortminus s,n+1}^{-}(\bm{q}_{\rm res})\lambda^{+}_{n+1,n+1+s}(\bm{0})}{E^{s}_{n+1+s}+s\hbar\omega_{c}-E_{n+1}^{-s}}\right|^{2}$
$\times N^{\rm el}_{n,s}\left(1{-}N^{\rm el}_{n+2,s}\right)\delta(\hbar\omega_{\rm res}{+}E_{n}^{s}{-}\hbar\omega_{c}{-}E_{n+2}^{s}),$ (7)
$r{=}\frac{2\pi}{\hbar}\!\!\sum_{\begin{subarray}{c}\bm{p}_{n},\bm{q}_{\rm res}\\ {s=\pm 1}\end{subarray}}\!\left|\frac{\lambda^{-}_{n+1+s,n+1}(\bm{0})\lambda_{n+1,n+1\shortminus s}^{+}(\bm{q}_{\rm res})}{E^{s}_{n+1+s}+s\hbar\omega_{c}-E_{n+1}^{-s}}\right|^{2}$
$\times N^{\rm el}_{n+2,s}\left(1{-}N^{\rm el}_{n,s}\right)\delta(\hbar\omega_{\rm res}{+}E_{n}^{s}{-}\hbar\omega_{c}{-}E_{n+2}^{s}).$ (8)
The 3D momentum of the pumping photon is denoted by $\bm{q}_{\rm res}$, electron level occupation is introduced as
@@
The diagrams in Fig. 5 describe photon absorption by the 2DEG with phonon emission.
@@
The rates of spontaneous and stimulated processes for a photonic mode of 2D momentum $\bm{q}$ are given by the following expressions for $\varkappa_{\rm loss}^{\prime}(\bm{q})$ and $\varkappa_{\rm loss}(\bm{q})$ respectively:
$\varkappa_{\rm loss}^{\prime}(\bm{q}){=}\frac{2\pi}{\hbar}\!\!\sum_{\begin{subarray}{c}p_{x},\bm{k}\\ {n,m,\zeta}\end{subarray}}\!\left|\!\frac{\lambda^{-}_{n,n\shortminus 1}(\bm{q})\kappa_{n\shortminus 1,m}(\bm{k},\zeta)}{E^{+}_{n}{+}\hbar\omega_{\bm{q}}{-}E_{n\shortminus 1}^{-}}{+}\frac{\kappa_{n,m+1}(\bm{k},\zeta)\lambda^{-}_{m+1,m}(\bm{q})}{E^{-}_{m}{-}\hbar\omega_{\bm{q}}{-}E_{m+1}^{+}}\!\right|^{2}\!\!\frac{N^{\rm el}_{m,\shortminus 1}\left({1{-}N^{\rm el}_{n,+1}}\right)}{e^{E_{\rm ph}(\bm{k},\zeta)/T}-1}\delta(\hbar\omega_{\bm{q}}{+}E_{n}^{+}{-}E_{m}^{-}{-}E_{\rm ph}({\bm{k}},\zeta)).$ (9)
$\varkappa_{\rm loss}(\bm{q})$ ${=}\frac{2\pi}{\hbar}\!\!\sum_{\begin{subarray}{c}p_{x},\bm{k}\\ {n,m,\zeta}\end{subarray}}\!\left|\!\frac{\lambda^{-}_{n,n\shortminus 1}(\bm{q})\kappa_{n\shortminus 1,m}(\bm{k},\zeta)}{E^{+}_{n}{+}\hbar\omega_{\bm{q}}{-}E_{n\shortminus 1}^{-}}{+}\frac{\kappa_{n,m+1}(\bm{k},\zeta)\lambda^{-}_{m+1,m}(\bm{q})}{E^{-}_{m}{-}\hbar\omega_{\bm{q}}{-}E_{m+1}^{+}}\!\right|^{2}\!\!\!\delta(\hbar\omega_{\bm{q}}{+}E_{n}^{+}{-}E_{m}^{-}{-}E_{\rm ph}({\bm{k}},\zeta))$
${\times}\left(\frac{N^{\rm el}_{n,+1}{-}N^{\rm el}_{m,\shortminus 1}}{e^{E_{\rm ph}(\bm{k},\zeta)/T}-1}+N^{\rm el}_{n,+1}\left(1{-}N^{\rm el}_{m,\shortminus 1}\right)\right)$ (10)
We further suppress the momentum argument for the condensate mode, thus $\varkappa_{\rm loss}=\varkappa_{\rm loss}(\bm{q}=0)$, $\varkappa^{\prime}_{\rm loss}=\varkappa^{\prime}_{\rm loss}(\bm{q}=0)$.
@@
for optical ones. The quantities $\Delta n_{a/{\rm opt}}$ denote the number of Landau levels between the scattering states of an electron.